OpenStudy (anonymous):

Anyone able to help me find the solution for f(n) in terms of f(0) for the recurrence relation: \[f\left(0 \right) = a\]\[f\left(n \right) = f\left(n-1\right)^2+b\]?

OpenStudy (amistre64):

write it out for starters, see if theres an apparent pattern to spot

OpenStudy (anonymous):

Tried it. It gets rather complicated.

OpenStudy (amistre64):

f0 = a f1= a^2 + b f2=(a^2+b)^2 + b f3=((a^2+b)^2 + b)^2 + b it does doesnt it lol

OpenStudy (amistre64):

it might be one of those; homogenous part and nonhomogenous part type thing

OpenStudy (anonymous):

I couldn't find anything on this type though.. I'll carry on looking.

OpenStudy (anonymous):

Oops, Mathworld seems to have the exact form. Hopefully it'll help me..

OpenStudy (amistre64):

good luck :)

OpenStudy (jamesj):

*bookmark. I'm curious to see the answer

OpenStudy (amistre64):

an=(an-1)^2; a0=n a0=n a1=n^2 a2=n^4 a3=n^8 a4=n^16 an=n^(2^n) would be the homogenous part; then we would need to construct that particular part.

OpenStudy (amistre64):

spose the "n" choice for the unknown was a bad option :)

OpenStudy (amistre64):

\[f(n) = a^{2^n}\] might look more better

OpenStudy (anonymous):

Right, after searching a little bit I found out that it's called a Quadratic Map. (See http://mathworld.wolfram.com/QuadraticMap.html) Unfortunately: "The well-known recurrence \[x_{n+1}={x_n}^2+c\]that is often called "the" quadratic map is not in general solvable in closed form." At least it's been cleared up. Had no idea that it was pretty much the Mandelbrot set.. Thanks for the help, and sorry to waste your time.